OpenStudy (anonymous):

My last question yay!!! (3-2i)(3-2i) I got 13??? Maybe im wrong

OpenStudy (luigi0210):

Yea, you're suppose to have an i term somewhere still.

OpenStudy (anonymous):

well you can always check by expanding: (3-2i)(3-2i)=9-6i-6i+4i^2 = 9+4-12i = 13-12i. dont forget your i terms!

OpenStudy (aayushi.somani):

yeah me too getting the same as @LifeEngineer

OpenStudy (aayushi.somani):

@sahmad u forget to add i to ur terms

OpenStudy (anonymous):

I got it thanks so much!!! I see what I did :) off to bed finally

OpenStudy (nurali):

@LifeEngineer Ur answer is wrong (3-2i)(3-2i) Multiply 3*3-3*2i-2i*3+2i*2i 9-6i-6i+4i^2 9-12i+4i^2 i^2=-1 9-12i+4(-1) 9-12i-4 5-12i

OpenStudy (anonymous):

@Nurali Right you are! Forgot to carry my negative on the i^2... Its been a while since I've done these types of problems